# MLP-SLAM: Multilayer Perceptron-Based Simultaneous Localization and Mapping With a Dynamic and Static Object Discriminator

### v1

@@ IV-B Coarse Estimation Module
Stereo images are used for ORB feature extraction initially. The left image is utilized for object detection and tracking using YOLOv7 [23] and Deep Sort [24]. In comparison to CFP-SLAM, which uses an extended Kalman filter (EKF) and Hungarian algorithm to compensate for missed object detection, we employ Deep Sort for a more effective implementation of the same function. Leveraging class information from the object detection model, we utilize the Depth Filter Module to treat background features as static points and combine them with all static feature points outside of the bounding boxes to estimate the camera pose by minimizing the re-projection error as in equation (1).
@@ IV-C Multilayer Perceptron Module
Before deploying the MLP model to classify dynamic and static feature points, a sufficient dataset is needed for training. Therefore, a dataset containing dynamic and static feature points is created. Details of the dataset collection process are provided in Section IV-E. Once an adequate dataset is available, the structure of the MLP model needs to be designed. The MLP model consists of three parts: the input layer, the hidden layer, and the output layer. The input layer’s number of nodes corresponds to the dimension of the input feature, which is three in this case: the image intensity error $e_{I}$, the epipolar error $e_{D}$, and the re-projection error $e_{Re}$. Previous work has often overlooked some of these features, leading to incomplete information. After setting up the input layer, our attention shifts to the hidden layer, which is crucial for the MLP model’s classification performance. In this study, a two-layer hidden layer with ten nodes each is constructed. The final step is to create the output layer. In our case, an output layer with two nodes that denotes the dynamic and static classes is utilized.
With a constructed MLP model and sufficient data, training is essential for successful classification of dynamic and static feature points. The first step involves applying zero mean normalization to the datasets. Although min-max normalization is attempted to scale data between $[0,1]$, zero mean normalization proved more effective based on our experiments. Following normalization, the dataset is divided into training and testing sets in an 8:2 ratio, with ten percent of the training data allocated for validation during the training process. Validation is crucial for assessing the model’s learning status, identifying over-fitting, under-fitting, or optimal learning. Subsequently, a loss function is designed, with the choice being the Binary Cross-Entropy (BCE) loss, which is a classic loss function for classification problems. Before training, hyperparameters and optimizer settings are configured as detailed in Section V-A. The optimal MLP model training process is presented in Figure 3, showing a consistent decrease in both training loss and validation losses indicated successful learning. The Accuracy metric, which calculates the ratio between the number of correct classification cases and the number of total cases, is shown to be increasing.
The MLP prediction module begins by generating input feature $v_{i}$ for each feature point $x_{i}$ that is located in the bounding box and is not filtered by Depth Filter Module, then normalizing the input feature $v_{i}$ through zero mean normalization to eliminate dimension effects, using the same parameters as the dataset normalization. Finally, all input features are put into the model in packs of a predetermined size and are classified as static or dynamic.
@@ IV-D Fine Estimation
IV-E Data Collection
For an extended period, there has been a lack of publicly available datasets for directly evaluating dynamic and static feature points classification methods. Currently, all methods are assessed indirectly through SLAM metrics, which can lead to inaccuracies and instability as it overlooks the randomness inherent in the SLAM system. For bridging this gap, we build a publicly available datasets which contains over 50,000 dynamic and static feature points from sequences 00 and 01 of KITTI tracking datasets. Each feature points is saved as following format, $u_{1}$, $v_{1}$, $z_{1}$, $id_{1}$, $u_{2}$, $v_{2}$, $id_{2}$, $class$, $e_{I}$, $e_{Re}$, $e_{D}$. The ($u_{1}$, $v_{1}$), ($u_{2}$, $v_{2}$) stands for coordinates of feature point $x_{1}$ and $x_{2}$ in image frame respectively. And $z1$ means the depth of $x_{1}$, which is computed through the same way as in ORB-SLAM2. The $id_{1}$ and $id_{2}$ represent whether the image frames belong to $x_{1}$ or $x_{2}$. The $class$ is binary, either $dynamic$ or $static$. The most important properties are $e_{I}$, $e{Re}$ and $e_{D}$, and they represent image intensity error, re-projection error and epipolar error respectively. We use ground truth pose of each frame to generated them. Details of how they are computed is represented in formulas (3), (4) and (5).
@@ V-A Experimental Settings
Parameter $\eta$ in Depth Filter Module is set to $1.2$. Parameters related to the training part of the MLP model are as follows: the epoch is 150, the batch size is 8, the learning rate is 0.1, the learning rate scheduler is MultiStepLR [25] with $\gamma$ set to 0.1. The optimizer we choose is Adam [26]. We converted the TensorRT [27] version of the MLP model with a fixed input shape of (200, 3).
@@ V-B Classification Evaluation
As previously stated, we have compiled a dataset for classifying dynamic and static feature points, consisting of over 50,000 feature points. This dataset enables the direct evaluation of various dynamic and static feature classification methods, eliminating the need for indirect comparisons based on metrics from the SLAM field. Direct evaluation is advantageous as it is less influenced by other factors within the SLAM system, resulting in more reliable and accurate experimental outcomes.
Our experiment involves the assessment of various methods for feature discrimination, including our MLP based dynamic and static feature discriminator, Support Vector Machine (SVM), a method akin to PointSLOT, a method similar to CFP-SLAM, and a method focusing solely on epipolar error. SVM is a well-established classification technique in machine learning, widely used in text and hypertext categorization, image classification, and satellite data classification. In this study, SVM from the Scikit-learn library [28] is directly employed, trained, and evaluated on the dataset. Additionally, two classification methods for dynamic and static feature points, similar to PointSLOT and CFP-SLAM, are implemented in Python due to the unavailability of their source code. The PointSLOT solely focuses on re-projection error, while the CFP-SLAM considered epipolar error in addition. Furthermore, a discrimination method concentrating solely on epipolar error is also implemented. The evaluation of these methods is based on the metrics of Accuracy and F1-score. Accuracy reflects the model’s predictive correctness, with true positive (TP), false positive (FP), true negative (TN), and false negative (FN) used to denote classification outcomes.. And the Accuracy is described as following:
@@
[…]
@@
The experimental results are presented in Table I. According to Table I, we can see that our MLP-based dynamic and static feature points discriminator achieves superior performance in terms of both Accuracy and F1-Score compared to other methods. Additionally, the support-vector machine (SVM) outperforms other methods except for our MLP-based method. Both MLP and SVM belong to machine learning methods, indicating that machine learning-based methods are more suitable for this dynamic and static feature classification problem compared to traditional methods that require manual parameter tuning. The significant difference between the method that only considers epipolar error and the method similar to PointSLOT which only considers re-projection error, suggests that re-projection error is a more critical indicator than epipolar error for this problem.
@@
Sequences ORB-SLAM2 DynaSLAM DynaSLAM II VDO-SLAM us
@@ V-C Visual Odometry
KITTI tracking dataset is a publicly available collection of stereo visual odometry datasets. It comprises 11 training sequences and 10 testing sequences featuring urban and road scenes captured from a car’s perspective, including moving vehicles, along with corresponding GPS data.
@@
We conduct a comprehensive comparison between our MLP-based SLAM system and other SLAM systems in this section. Table II presents a detailed comparison result. In Table II, we provide an absolute translation error (ATE) comparison with ORB-SLAM2, DynaSLAM, DynaSLAM II, and a modified version of VDO-SLAM. ORB-SLAM2 does not specifically address dynamic situations but serves as the foundation of our SLAM system. DynaSLAM also builds upon ORB-SLAM2 with the ability to detect dynamic objects, but it discards features located within the bounding box of potential dynamic objects directly, leading to information loss. DynaSLAM II is the current state-of-the-art method, aiming to enhance camera pose estimation precision by optimizing the Bundle Adjustment (BA) factor graph, which includes object poses and constant velocity constraints. The modified version of VDO-SLAM [29] is a re-implemented version; the original version can only run on datasets provided by the authors. The modified VDO-SLAM requires acquiring depth images through ZoeDepth [30], optical flow through VideoFlow offline, and applying YOLOv8 [31] to detect dynamic objects. We tested all SLAM systems on our experimental devices except DynaSLAM II due to the unavailability of its source code. Therefore, we directly used the experimental results from its original paper.
@@
According to experimental results in Table II, it can be seen that our SLAM system outperforms other SLAM system in most of sequences of Kitti tracking datasets. More specifically, our SLAM system outperforms VDO-SLAM-modified version in all sequences. Besides, compared to our base ORB-SLAM2, we achieve better performance in the sequences 00, 01, 03, 05, 07, 08, 09 and 10. The sequence 01 is collected from a car driving on the high way, where all cars that show up at this sequence are dynamic. Our SLAM system works better than ORB-SLAM2 in this sequence, which indicates that our SLAM system has the ability to address dynamic objects in the surrounding environment correctly. Moreover, compared to DynaSLAM, our SLAM system achieves higher precision on sequences except for sequences 04 and 06, and the gap on the sequence 04 and 06 are within 0.1. Compared to DynaSLAM II, our SLAM system only performs worse on the sequences 01, 02, 06 and 08 out of 11 sequences. However, our SLAM system achieved lowest average absolute translation error (ATE). It indicates that our SLAM system achieves the highest average precision over all sequences of Kitti tracking dataset. It is worth mentioning that the sequence 00 is rich of static cars that are parked on the roadside, our SLAM system achieves the best performance in this sequence compared to other methods, which indicates that our MLP model is capable to identifying real dynamic objects.
TABLE III: Latency Of Each Module
Module ORB-SLAM2 (stereo) us
@@ V-D Time Analysis
Low latency is one of the desired property of the SLAM system. Most of onboard devices are not capable of executing tasks with high computation cost. We lists the average computational time of each module in the tracking thread of our MLP based stereo SLAM system in Table III. According to experimental results from Table II and Table III, our MLP-based SLAM system achieves higher precision with only 17.50ms/frame extra computation cost. Moreover, we made a comprehensive latency comparison with other SLAM systems. The mean values of the highest latency and the lowest latency from the original papers of DynaSLAM II and PointSLOT are presented in Table IV. For all other methods in Table IV, average latency value in Kitti tracking sequences from 00 to 10 are presented. It is worth mentioning that even though the latency of DynaSLAM II does not include the time consumption of 2d and 3d object detection models, the execution time of our methods is still 17.34ms faster. Combining Table II and Table IV, we can conclude that our MLP-based SLAM system achieves the lowest average absolute translation error (ATE) and the fastest speed compared to other dynamic SLAM systems.
TABLE IV: Latency Anaylysis
@@
us Stereo 62.81 ms
@@ VI CONCLUSIONS
In this letter, we propose a MLP-based real-time stereo SLAM system to address SLAM system degradation in dynamic environments. Currently, there is no publicly available dataset for directly evaluating the effectiveness of dynamic and static feature classification methods. Previous assessments have relied on indirect metrics, which are prone to inaccuracy due to the randomness in the Random Sample Consensus (RANSAC) process in camera pose estimation within the SLAM systems. To bridge this gap, we have created a publicly available dataset containing over 50,000 feature points. Our proposed MLP-based dynamic and static feature discriminator is the most effective method compared to other methods on our collected datasets. Additionally, our MLP-based real-time stereo SLAM system achieves the highest average precision on the KITTI tracking dataset and the fastest speed compared to other dynamic SLAM systems.

### v2

@@ IV-B Coarse Estimation Module
Stereo images are used for ORB feature extraction initially. The left image is utilized for object detection and tracking using YOLOv7 [wang2023yolov7] and Deep Sort [du2023strongsort]. In comparison to CFP-SLAM, which uses an extended Kalman filter (EKF) and Hungarian algorithm to compensate for missed object detection, we employ Deep Sort for a more effective implementation. Leveraging class information from the object detection model, we utilize the DFM to treat background features as static points and combine them with all static feature points outside of the bounding boxes to estimate the camera pose by minimizing the re-projection error as in equation (1).
@@ IV-C Multilayer Perceptron Module
Before deploying the MLP model for classifying feature points, a sufficient training dataset of dynamic and static points is collected. Details of the dataset collection process are provided in Section IV-E. Once an adequate dataset is available, the structure of the MLP model needs to be designed. The MLP model consists of three parts: the input layer, the hidden layer, and the output layer. The input layer’s number of nodes corresponds to the dimension of the input feature, which is three in this case: the image intensity error $e_{I}$, the epipolar error $e_{D}$, and the re-projection error $e_{Re}$. Previous work like[CFP-SLAM] and [PointSLOT] has often overlooked some of these features, leading to incomplete information.
Training is crucial for classifying dynamic and static feature points. The process begins with zero-mean normalization of the datasets, followed by dividing the dataset into training, testing, and validation sets. Before training, hyperparameters and optimizer settings are configured as detailed in Section V-A. After lots of experiments, the optimal MLP model training process is presented in Figure 4, showing a consistent decrease in both training loss and validation losses indicated successful learning. The Accuracy metric, which calculates the ratio between the number of correct classification cases and the number of total cases, is shown to be increasing.
The MLP prediction module begins by generating input feature $v_{i}$ for each feature point $x_{i}$ that is located in the bounding box and is not filtered by DFM, then normalizing the input feature $v_{i}$ through zero mean normalization to eliminate dimension effects, using the same parameters as the dataset normalization. Finally, all input features are put into the model in batches of a predetermined size and are classified as static or dynamic.
@@ IV-D Fine Estimation
IV-E Data Collection And Data Split
At present, all methods for classifying dynamic and static feature points are evaluated indirectly through SLAM metrics like absolute translation error (ATE), relative translation error (RTE) and relative rotation error (RRE), primarily due to the absence of a dedicated classification dataset. This reliance on indirect assessment may result in inaccuracies and instability, as it fails to account for the inherent randomness present in SLAM systems. By establishing a comprehensive classification dataset, it becomes possible to evaluate dynamic and static feature point classification methods directly, thereby enhancing the reliability of experimental results. For bridging this gap, we build a publicly available datasets which contains over 50,000 dynamic and static feature points from sequences 00 and 01 of KITTI Odometry datasets. All those feature points manually labeled, and the ratio of dynamic and static feature points in collected dataset is 1:1. Each feature points is saved as following format, $u_{1}$, $v_{1}$, $z_{1}$, $id_{1}$, $u_{2}$, $v_{2}$, $id_{2}$, classclass, $e_{I}$, $e_{Re}$, $e_{D}$. The ($u_{1}$, $v_{1}$), ($u_{2}$, $v_{2}$) stands for coordinates of feature point $x_{1}$ and $x_{2}$ in image frame respectively. And $z_{1}$ means the depth of $x_{1}$, which is computed through the same way as in ORB-SLAM2. The $id_{1}$ and $id_{2}$ represent whether the image frames belong to $x_{1}$ or $x_{2}$. The classclass is binary, either dynamicdynamic or staticstatic. The most important properties are $e_{I}$, $e_{Re}$ and $e_{D}$, and they represent image intensity error, re-projection error and epipolar error respectively. We use ground truth pose of each frame to generated them. Details of how they are computed is represented in Eq. (3), (4) and (5).
A proper dataset split is crucial for training supervised machine learning models. We shuffle the dataset of over 50,000 feature points at first, then allocate 70% as training set, 20% as testing set, and 10% as validation set. Besides, We maintain a 1:1 ratio of static to dynamic feature points in all sets to address data imbalance.
@@ V-A Experimental Settings
Parameter $\eta$ in the DFM is set to $1.2$. Parameters related to the training part of the MLP model are as follows: the epoch is 150, the batch size is 8, the learning rate is 0.1, the learning rate scheduler is MultiStepLR [paszke2019pytorch] with $\gamma$ set to 0.1. The optimizer we choose is Adam [kingma2014adam]. We converted the TensorRT [tensorRT] version of the MLP model with a fixed input shape of (200, 3).
@@ V-B Classification Evaluation
We have compiled a dataset to classify dynamic and static feature points, consisting of over 50,000 feature points. This dataset enables the direct evaluation of various dynamic and static feature classification methods, eliminating the need for indirect comparisons based on metrics from the SLAM field. Direct evaluation is advantageous as it is less influenced by other factors within the SLAM system, resulting in more reliable and accurate experimental outcomes.
Our experiment involves the assessment of various methods for feature discrimination, including our MLP based dynamic and static feature discriminator, Support Vector Machine (SVM), a method akin to PointSLOT, a method similar to CFP-SLAM, and a method focusing solely on epipolar error. SVM is a well-established classification technique in machine learning, widely used in text and hypertext categorization, image classification, and satellite data classification. In this study, SVM from the Scikit-learn library [pedregosa2011scikit] is directly employed, trained, and evaluated on the dataset. Additionally, two classification methods for dynamic and static feature points, similar to PointSLOT and CFP-SLAM, are implemented in Python due to the unavailability of their source code. The MLP based method and SVM are trained on the same training set, and all five methods are tested on the same testing set. The training set, testing set and validation set are obtained through the data split method in Section IV-E. Experimental results presented in Table I are the best performance in the testing set of collected dataset by fine-tuning the hyper-parameters of the methods. The PointSLOT solely focuses on re-projection error, while the CFP-SLAM considered epipolar error in addition. Furthermore, a discrimination method concentrating solely on epipolar error is also implemented. The evaluation of these methods is based on the metrics of Accuracy and F1-score. Accuracy reflects the model’s predictive correctness, with true positive (TP), false positive (FP), true negative (TN), and false negative (FN) used to denote classification outcomes.. And the Accuracy is described as following:
@@
$\text{F1-Score}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (7)
@@
The experimental results are presented in Table I. According to it, we can see that our MLP-based method achieves superior performance in terms of both Accuracy and F1-Score compared to other methods. Additionally, the support-vector machine (SVM) outperforms other methods except for our MLP-based method, indicating that machine learning-based methods are more suitable for this dynamic and static feature classification problem compared to traditional methods that require manual parameter tuning. The significant difference between the method that only considers epipolar error and the method similar to PointSLOT which only considers re-projection error, suggests that re-projection error is a more critical indicator than epipolar error for this problem.
@@
Sequences ORB-SLAM2 DynaSLAM DynaSLAM II VDO-SLAM Ours
@@
TABLE III: The ATE Comparison on KITTI360 Datasets
Sequences ORB-SLAM2 VDO-SLAM Ours
00 1.46 6.92 1.36
01 2.83 10.73 2.04
02 6.82 6.17
03 4.58 4.11
average 3.92 8.83 3.42
@@ V-C Visual Odometry
KITTI Odometry dataset is a publicly available collection of stereo visual odometry data, with 11 training and 10 testing sequences featuring urban and road scenes from a car’s perspective, including moving vehicles and GPS data. KITTI360 enhances this dataset with better localization and richer semantic information.
@@
We conduct a comprehensive comparison between our MLP-based SLAM system and other SLAM systems in this section. A presentation video of our experiment can be found in https://www.youtube.com/watch?v=sOKNc0-ZH2Q. In Table II, we provide an absolute translation error (ATE) comparison with ORB-SLAM2, DynaSLAM, DynaSLAM II, and a modified version of VDO-SLAM on the KITTI Odometry dataset. It is important to note that we deleted portions of the image frames in sequences 00 and 01 of KITTI Odometry dataset which were used for collecting the training dataset for the MLP model. Moreover, comparison with ORB-SLAM2, DynaSLAM, modified VDO-SLAM on the KITTI360 dataset is presented in Table III. The MLP model for distinguishing dynamic and static feature points was trained only on the KITTI Odometry dataset and did not use any possible information from the KITTI360 dataset. ORB-SLAM2 does not specifically address dynamic situations but serves as the foundation of our SLAM system. DynaSLAM also builds upon ORB-SLAM2 with the ability to detect dynamic objects, but it discards features located within the bounding box of potential dynamic objects directly, leading to information loss. DynaSLAM II is the current state-of-the-art method, aiming to enhance camera pose estimation precision by optimizing the Bundle Adjustment (BA) factor graph, which includes object poses and constant velocity constraints. The modified version of VDO-SLAM [vdo-slam-modified] is a re-implemented version; the original version can only run on datasets provided by the authors. The modified VDO-SLAM requires acquiring depth images through ZoeDepth [bhat2023zoedepth], optical flow through VideoFlow offline, and applying YOLOv8 [Jocher_Ultralytics_YOLO_2023] to detect dynamic objects. We tested all SLAM systems on our experimental devices except DynaSLAM II due to the unavailability of its source code. Therefore, we directly used the experimental results from its original paper.
@@
According to experimental results in Table II and Table III, it can be seen that our SLAM system outperforms other SLAM system in most of sequences of KITTI Odometry and KITTI360 dataset. More specifically, our SLAM system outperforms VDO-SLAM-modified version in all sequences. Besides, compared to our base ORB-SLAM2, we achieve better performance in the sequences 00, 01, 03, 05, 07, 08, 09 and 10 of KITTI Odometry dataset and in all sequences of KITTI360 dataset. The sequence 01 was collected from a car on the highway, featuring dynamic vehicles. Our SLAM system surpasses ORB-SLAM2 in this sequence, proving its ability to effectively manage dynamic objects in the environment. Moreover, compared to DynaSLAM, our SLAM system achieves higher precision on sequences except for sequences 04 and 06 of KITTI Odometry dataset, and the gap on the sequence 04 and 06 are within 0.1. Compared to DynaSLAM II, our SLAM system only performs worse on the sequences 01, 02, 06 and 08 out of 11 sequences. Our SLAM system recorded the lowest average absolute translation error (ATE) on the KITTI Odometry and KITTI360 datasets, demonstrating the highest average precision across all sequences. It is worth mentioning that the sequence 00 is rich of static cars that are parked on the roadside, our SLAM system achieves the best performance in this sequence compared to other methods, which indicates that our MLP model is capable to identifying real dynamic objects. Although experiments are conducted solely using a stereo sensor, our proposed method can be extended to accommodate monocular or RGB-D cameras.
TABLE IV: Latency Of Each Module
Module ORB-SLAM2 (stereo) Ours
@@ V-D Time Analysis
Low latency is one of the desired property of the SLAM system. We lists the average computational time of each module in the tracking thread of our MLP based stereo SLAM system in Table IV. According to experimental results from Table II and Table IV, our MLP-based SLAM system achieves higher precision with only 17.50 ms/frame extra computation cost compared to ORB-SLAM2. Moreover, we made a comprehensive latency comparison with other SLAM systems. The mean values of the highest latency and the lowest latency from the original papers of DynaSLAM II and PointSLOT are presented in Table V. For all other methods in Table V, average latency value in KITTI Odometry sequences from 00 to 10 are presented. It is worth mentioning that even though the latency of DynaSLAM II does not include the time consumption of 2d and 3d object detection models, the execution time of our methods is still 17.34 ms faster. Combining Table II, Table III and Table V, we can conclude that our MLP-based SLAM system achieves the lowest average absolute translation error (ATE) and the fastest speed compared to other dynamic SLAM systems.
TABLE V: Latency Anaylysis
@@
Ours Stereo 62.81 ms
V-E Ablation Study
We evaluated our MLP module’s effectiveness for dynamic feature point removal by comparing three models.Our SLAM system features just the object tracking module. Our SLAM system uses only the MLP module.And SLAM system features both modules and additional components stayed the same. The experimental results are presented in Table VI. It indicated that the SLAM system with MLP-only module outperforms the object tracking module, and their combination achieves the best performance.
TABLE VI: The ATE Comparison on KITTI Datasets of Abalation Study
Sequences object tracking only MLP only both
00 2.53 1.37 0.79
01 10.71 6.79 4.53
02 5.73 4.01 3.48
03 0.42 0.34 0.34
04 0.18 0.17 0.24
05 0.38 0.35 0.33
06 1.03 0.80 0.88
07 0.49 0.43 0.41
08 3.36 2.82 3.18
09 3.43 3.31 0.90
10 1.16 1.23 1.27
average 2.67 1.97 1.48
@@ VI CONCLUSIONS
In this letter, we propose a MLP-based real-time stereo SLAM system to address SLAM system degradation in dynamic environments. Currently, there is no publicly available dataset for directly evaluating the effectiveness of dynamic and static feature classification methods. Previous assessments have relied on indirect metrics, which are prone to inaccuracy due to the randomness in the Random Sample Consensus (RANSAC) process in camera pose estimation within the SLAM systems. To bridge this gap, we have created a publicly available dataset containing over 50,000 feature points. Our proposed MLP-based dynamic and static feature discriminator is the most effective method compared to other methods on our collected datasets. Additionally, our MLP-based real-time stereo SLAM system achieves the highest average precision on the KITTI Odometry dataset and the fastest speed compared to other dynamic SLAM systems.